Exercise 16-26

Headland Company’s net income for 2020 is $49,300. The only potentially dilutive securities outstanding were 1,200 options issued during 2019, each exercisable for one share at $6. None has been exercised, and 9,300 shares of common were outstanding during 2020. The average market price of Headland’s stock during 2020 was $18.

(a) Compute diluted earnings per share. (Round answer to 2 decimal places, e.g. $2.55.)

Diluted earnings per share $


(b) Assume the same facts as those assumed for part (a), except that the 1,200 options were issued on October 1, 2020 (rather than in 2019). The average market price during the last 3 months of 2020 was $18. (Round answer to 2 decimal places, e.g. $2.55.)

Diluted earnings per share $
Answer #1

Ans:

a)

Basic EPS of the firm = 49,300 / 9,300

$      5.30

Assuming that the options were exercised at the beginning of the year

Incremental number of shares = (Market price - Option price / Market price) x Number of options=(18-6/18)*1200

800

Diluted EPS = 49300 + 0 / 9300+800

$      4.88

b)

b)

Diluted EPS = 49300 + 0 / 9300+(800 x 3/12)

$      5.19


Hope this helped ! Let me know in case of any queries.